# MyFramelinecolor=black, outerlinewidth=2pt, roundcorner=20pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=codeBackground

### v3

@@ 5.1 A (Trivial) PoUW for Any Function
We show that a PoUW for any function $f$ with an associated Alg running in time $t(n)$ on instances of size $n$ exists, albeit this scheme is not optimal. Specifically, there is a gap between the complexity of $P$ and of the associated algorithm, and also there is a $P^{\star}$ that has significantly less complexity than $P$. At a very high level, the scheme is doing a computation plus proof of work, independently. More details follow.
  • $\mathsf{Solve}$: on input $\sigma$ and $\mathsf{x}$, the prover solves the task by running Alg on $\mathsf{x}$ to get $\mathsf{y}$. Also, independently, it does a proof of work to convince the verifier that it spent enough time. Specifically, for a constant $c$ or perhaps slightly super-constant (see discussion below), it chooses $T=c\cdot t(n)$ nonces $\omega_{1},\ldots,\omega_{T}$ for the random oracle, and computes $z_{i}=\mathcal{O}(\sigma,\omega_{i})$. It then sets $\pi$ to be the $\omega_{i}$ for which $z_{i}$ starts with the most 00’s.
    The prover’s running time is thus $\alpha(t(n))=(1+c)\cdot t(n)$.
  • $\mathsf{Verify}$: on input $\sigma$ and $\omega$, the verifier outputs $1$ if and only if $\mathcal{O}(\sigma,\omega)$ has $\log(t(n))$ leading 00’s.
    The verifier’s running time is thus $\beta(t(n))=\tilde{O}(1)$.
@@
The above PoUW protocol satisfies $(1-\epsilon)$-completeness and hardness (see Definition 5.1) with $\epsilon=e^{-c}$.
@@
Since $\mathcal{O}$ is a random oracle, its output on previously unqueried points is uniformly distributed in $\{0,1\}^{\lambda}$. Furthermore, the probability of success for a single query is $p=1/t(n)$. The probability that all $T$ attempts fail, is $(1-p)^{c\cdot t(n)}\leq e^{-c}$. Thus, $(1-\epsilon)$-correctness of proof holds. Hardness follows directly, again, from the random oracle property. A method for “guessing” a nonce that leads to an output that has $\log T$ leading 0s amounts to “guessing” the output of a random oracle without even making the appropriate query. ∎
The protocol is not an optimal PoUW. Indeed, either $P$ has linear overhead over the complexity of Alg, or an adversarial $P^{\star}$ can gain linear advantage in running time over $P$. Details follow. Consider the following malicious prover $P^{\star}$: $P^{\star}$ does not multiply any two matrices, and instead only finds the good nonce out of $T$ arbitrary ones. Clearly, this $P^{\star}$ has exactly the same probability of convincing $V$ as $P$, but it is more efficient. Concretely, in this protocol protocol $t_{P^{\star}}(n)=c\cdot t(n)$ which is a factor $(1+1/c)$ faster than $t_{P}(n)$. It is possible to make the advantage of $P^{\star}$ sub-linear by setting $c=\omega(1)$, but this takes the other issue to the extreme: the honest prover $P$’s complexity is a multiplicative factor $(1+c)$ larger than the complexity of Alg.
@@ 6.1 The Canonical MatMul Algorithm
{mdframed}
[ backgroundcolor=codeBackground, linecolor=black, linewidth=0.5pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]Canonical Matrix Multiplication Algorithm
@@
  1. 1.
    Initialize an all-00 matrix $C^{(0)}\in[-q,q]^{n\times m}$.
  2. 2.
    Partition $A$, $B$, and $C$ into blocks of size $r\times r$. The $(i,j)$th block of a matrix $X\in\{A,B,C\}$ is denoted $X_{i,j}$.
  3. 3.
    For each $i\in[n/r]$, $j\in[m/r]$, and $\ell\in[k/r]$, compute
    $C^{(\ell)}_{i,j}:=C^{(\ell-1)}_{i,j}+A_{i,\ell}\cdot B_{\ell,j},$
    where $A_{i,\ell}\cdot B_{\ell,j}$ can be computed via any classical matrix multiplication algorithm.
  4. 4.
    Output $C^{(n)}$.
Algorithm 6.1
@@
The first idea is to use fast square matrix multiplication algorithms. Specifically, every intermediate value is obtained by adding to a previously calculated intermediate value the product of two $r\times r$ matrices. Since this can be done in time $r^{\omega}$ where $\omega$ is the exponent of matrix multiplication (currently standing at $\omega=2.371552$ [VXXZ24]), we obtain an algorithm with overall complexity $O(nmk/r^{3-\omega})$, beating the naive one.
A somewhat more efficient approach in many cases is to utilize fast rectangular matrix multiplication algorithms. Specifically, consider the matrix induced by the values $C_{i,j}^{\ell}$ for a fixed $\ell$. Let us focus on $\ell=1$ for concreteness. This matrix is, in fact, the product of the left-most column-strip of $A$ (when we view the matrix as having $r\times r$ atomic blocks as elements) and the top-most row-strip of $B$ (essentially computing the outer product of these vectors). We shall denote $\omega_{m}$ the exponent of the best matrix multiplication algorithm for multiplying $n\times m$ and $m\times n$ matrices for $m<n$. We get an algorithm that runs in time $(n/r)\cdot n^{\omega_{r}}=n^{\omega_{r}+1}/r$, a term that could be $o(n^{3})$. As a concrete setting, the current record says that computing an $n\times n^{k}$ times an $n^{k}\times n$ matrix for all $k\leq 0.321334$ can be done in time $n^{2+o(1)}$ [Gal24, VXXZ24]. That is $\omega_{k}=2$ for all $k\leq 0.321334$. For $r=n^{0.3}$, this results with an algorithm that works in time $n^{2.7+o(1)}$ for computing all of the intermediates.
@@ 6.2 A Transcript Unpredictable Encoding Scheme
[…]
@@
[…]
@@
[…]
@@
$\epsilon$-transcript unpredictability: For every two $n\times n$ matrices $A,B$, and every $r$, every iterative numerical algorithm $\mathcal{A}$ that runs in time $o(t_{\textsc{MatMul}}(n))$ cannot compute the transcript of the product of the noisy matrices $A^{\prime},B^{\prime}$, except with probability $\epsilon$. That is,
$\Pr\left[\begin{aligned} &\forall i,j,k\in[n/r]\colon\\ &\quad C^{(k)}_{i,j}=Z^{(k)}_{i,j}\end{aligned}\,\middle|\,\begin{aligned} &A^{\prime},B^{\prime}=\textsc{Encode}_{r}(\sigma,A,B)\\ &\{C^{(k)}_{i,j}\}_{i,j,k\in[n/r]}=\mathsf{Tr}(\textsc{MatMul}_{r},A^{\prime},B^{\prime})\\ &\mathcal{\{}Z^{(k)}_{i,j}\}_{(i,j,k)\in[n/r]}=\mathcal{A}(A^{\prime},B^{\prime})\\ \end{aligned}\right]\leq\epsilon,$
@@ 6.3 From Transcript Unpredictability to a PoUW
{mdframed}
[ backgroundcolor=codeBackground, linecolor=black, linewidth=0.5pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]A PoUW for Matrix Multiplication with Parameter $r$
@@
  1. 1.
    Compute $A^{\prime},B^{\prime}=\textsc{Encode}_{r}(\sigma,A,B)$.
  2. 2.
    Computes $C^{\prime}=\textsc{MatMul}_{r}(A^{\prime},B^{\prime})$.
    Denote $z=\{C^{(k)}_{i,j}\}_{i,j,k\in[n/r]}$ the intermediate $r\times r$ matrices.
  3. 3.
    Compute $C:=\textsc{Decode}_{r}(\sigma,C^{\prime})$.
  4. 4.
    Output $(C,\pi)$ where $\pi=(A,B,z)$.
$\mathsf{Verify}(\sigma,\pi)$:
  1. 1.
    Parse $\pi=(A,B,z)$.
  2. 2.
    Recomputes from $\sigma,A,B$ the correct value of $z$, and output $1$ if and only if $z$ is correct.
Algorithm 6.2
@@
[…]
@@ 6.4 A First Instantiation
{mdframed}
[ backgroundcolor=codeBackground, linecolor=black, linewidth=0.5pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]First Implementation of Encode and Decode
$\textsc{Encode}_{r}(\sigma,A,B)$: $\quad\sslash\quad A,B\in{\mathbb{F}}_{q}^{n\times n}$
  1. 1.
    Parse $\sigma=E,F$, where $E,F\in{\mathbb{F}}_{q}^{n\times n}$.
  2. 2.
    Output $A^{\prime}=A+E$ and $B^{\prime}=B+F$.
$\textsc{Decode}_{r}(\sigma,C^{\prime})$: $\quad\sslash\quad C^{\prime}\in{\mathbb{F}}_{q}^{n\times n}$
  1. 1.
    Parse $\sigma=E,F$, where $E,F\in{\mathbb{F}}_{q}^{n\times n}$.
  2. 2.
    Compute $C^{\prime\prime}=A\cdot F+E\cdot(B+F)$ (using two invocations of $\textsc{MatMul}_{r}$ and a matrix addition).
  3. 3.
    Output $C=C^{\prime}-C^{\prime\prime}$.
Algorithm 6.3
@@
The obvious downside is that a malicious worker can skip the Decode operation altogether and thereby save 2 matrix multiplications. In other words, the honest worker does at least 3 times the work it should have done only for computing $\textsc{MatMul}_{r}$, i.e., the protocol is $\alpha$-efficient for $\alpha(t_{\textsc{MatMul}}(n))=3(t_{\textsc{MatMul}}(n)+o(1))$. On the other hand, there is a $\mathsf{Solve}^{*}$ that does essentially only $t_{\textsc{MatMul}}(n)$ work, concluding that the protocol is not an optimal PoUW.
@@ 6.5 A Second Instantiation
{mdframed}
[ backgroundcolor=codeBackground, linecolor=black, linewidth=0.5pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]Second Implementation of Encode and Decode
$\textsc{Encode}_{r}(\sigma,A,B)$: $\quad\sslash\quad\sigma\in\{0,1\}^{\lambda},\quad A,B\in{\mathbb{F}}_{q}^{n\times n}$
  1. 1.
    Parse $\sigma=E_{L},E_{R},F_{L},F_{R}$ as four matrices, where $E_{L},F_{L}\in{\mathbb{F}}_{q}^{n\times r}$ and $E_{R},F_{R}\in{\mathbb{F}}_{q}^{r\times n}$.
  2. 2.
    Compute $E=E_{L}\cdot E_{R}\text{ and }F=F_{L}\cdot F_{R}$ (invoking MatMul twice).
  3. 3.
    Output $A^{\prime}=A+E$ and $B^{\prime}=B+F$.
$\textsc{Decode}_{r}(\sigma,C^{\prime})$: $\quad\sslash\quad\sigma\in\{0,1\}^{\lambda},\quad C^{\prime}\in{\mathbb{F}}_{q}^{n\times n}$
  1. 1.
    Parse $\sigma=E_{L},E_{R},F_{L},F_{R}$ as four matrices, where $E_{L},F_{L}\in{\mathbb{F}}_{q}^{n\times r}$ and $E_{R},F_{R}\in{\mathbb{F}}_{q}^{r\times n}$.
  2. 2.
    Compute $C^{\prime\prime}=(A\cdot F_{L})\cdot F_{R}+E_{L}\cdot(E_{R}\cdot(B+F_{L}\cdot F_{R}))$ (invoking MatMul 5 times).
  3. 3.
    Output $C=C^{\prime}-C^{\prime\prime}$.
Algorithm 6.4
@@
Here, as opposed to all of our previous suggestions, the PoUW is optimal. Indeed, the honest solver $\mathsf{Solve}$ is doing exactly one product of two $n\times n$ matrices, as all of the products in Encode and Decode are products of $n\times n$ and $n\times r$ matrices, that cost $O(n^{2}\cdot r)$. Overall, by a direct analysis,
$\alpha(t_{\textsc{MatMul}}(n))=O(n^{2}+n^{2}\cdot r)+t_{\textsc{MatMul}}(n).$
Note that the last summand is the dominant term. By our assumption on the fastest way to compute the intermediate matrices is by using fast (rectangular) matrix multiplication as described in Section 6.1, there is no algorithm that runs in time $o(t_{\textsc{MatMul}}(n))$, making our protocol an optimal PoUW.

### v4

@@ 5.1 A (Trivial) PoUW for Any Function
We show that a PoUW for any function $f$ with an associated Alg running in time $t(n)$ on instances of size $n$ exists, albeit this scheme is not very useful since its associated $\alpha$ is large. Specifically, there is a gap between the complexity of $P$ and of the associated algorithm, and also there is a $P^{\star}$ that has significantly less complexity than $P$ and can generate “valid” $\pi$’s. At a very high level, the scheme is doing a computation plus proof of work, independently, as follows.
  • $\mathsf{Solve}$: on input $\sigma$ and $\mathsf{x}$, the prover solves the task by running Alg on $\mathsf{x}$ to get $\mathsf{y}$. Also, independently, it does a proof of work to convince the verifier that it spent enough time. Specifically, for a constant $c$ or perhaps slightly super-constant (see discussion below), it chooses $T=c\cdot t(n)$ nonces $\omega_{1},\ldots,\omega_{T}$ for the random oracle, and computes $z_{i}=\mathcal{O}(\sigma,\omega_{i})$. It then sets $\pi$ to be the $\omega_{i}$ for which $z_{i}$ starts with the most 0’s.
  • $\mathsf{Verify}$: on input $\sigma$ and $\omega$, the verifier outputs $1$ if and only if $\mathcal{O}(\sigma,\omega)$ has $\log(t(n))$ leading 0’s.
    The verifier’s running time is thus $\tilde{O}(1)$.
@@
The above PoUW protocol with overhead $\alpha=c$ and difficulty $\epsilon=1-e^{-c}$.
@@
Since $\mathcal{O}$ is a random oracle, its output on previously-unqueried points is uniformly distributed in $\{0,1\}^{\lambda}$. Overhead is $\alpha=c$ directly by construction. Furthermore, the probability of success for a single query is $p=1/t(n)$. The probability that all $T$ attempts fail, is $(1-p)^{c\cdot t(n)}\leq e^{-c}$, so completeness holds. Hardness follows from the random oracle property: A method for “guessing” a nonce that leads to an output that has $\log T$ leading 0s amounts to “guessing” the output of a random oracle without even making the appropriate query. ∎
The protocol is not useful in applications: either $P$ has linear overhead over the complexity of Alg, or an adversarial $P^{\star}$ can gain linear advantage in running time over $P$. Consider the following malicious prover $^{\star}: PP^{\star}$ does not multiply any two matrices, and instead only finds the good nonce out of $T$ arbitrary ones. Clearly, this $P^{\star}$ has exactly the same probability of convincing $V$ as $P$, but it is more efficient. Concretely, in this protocol protocol $t_{P^{\star}}(n)=c\cdot t(n)$ which is a factor $(1+1/c)$ faster than $t_{P}(n)$. It is possible to make the advantage of $P^{\star}$ sub-linear by setting $c=\omega(1)$, but this takes the other issue to the extreme: the honest prover $P$’s complexity is a multiplicative factor $(1+c)$ larger than the complexity of Alg.
@@ 6.1 The Canonical MatMul Algorithm
The first idea is to use fast square matrix multiplication algorithms. Specifically, every intermediate value is obtained by adding to a previously calculated intermediate value the product of two $r\times r$ matrices. Since this can be done in time $r^{\omega}$ where $\omega$ is the exponent of matrix multiplication (currently standing at $\omega=2.371552$ [DBLP:conf/soda/WilliamsXXZ24]), we obtain an algorithm with overall complexity $O(nmk/r^{3-\omega})$, beating the naive one.
A somewhat more efficient approach in many cases is to utilize fast rectangular matrix multiplication algorithms. Specifically, consider the matrix induced by the values $C_{i,j}^{\ell}$ for a fixed $\ell$. Let us focus on $\ell=1$ for concreteness. This matrix is, in fact, the product of the left-most column-strip of $A$ (when we view the matrix as having $r\times r$ atomic blocks as elements) and the top-most row-strip of $B$ (essentially computing the outer product of these vectors). We shall denote $\omega_{m}$ the exponent of the best matrix multiplication algorithm for multiplying $n\times m$ and $m\times n$ matrices for $m<n$. We get an algorithm that runs in time $(n/r)\cdot n^{\omega_{r}}=n^{\omega_{r}+1}/r$, a term that could be $o(n^{3})$. As a concrete setting, the current record says that computing an $n\times n^{k}$ times an $n^{k}\times n$ matrix for all $k\leq 0.321334$ can be done in time $n^{2+o(1)}$ [DBLP:conf/soda/Gall24, DBLP:conf/soda/WilliamsXXZ24]. That is $\omega_{k}=2$ for all $k\leq 0.321334$. For $r=n^{0.3}$, this results with an algorithm that works in time $n^{2.7+o(1)}$ for computing all of the intermediates.
@@ 6.2 A Transcript Unpredictable Encoding Scheme
Our template relies on a generic invertible operation we call a transcript unpredictable encoding scheme. The goal of the encoding procedure of the scheme is to inject noise and perturb the input matrices so that the transcript of the perturbed matrices has non-trivial entropy. The inverse operation allows to “peel off” the injected noise and recover the original output. Both operations are parametrized with $r\in[n]$, the same parameter that appears in $\textsc{MatMul}_{r}$. The scheme consists of two deterministic algorithms $(\textsc{Encode},\textsc{Decode})$ with the following syntax:
@@
We now formalize the properties of a transcript unpredictable encoding scheme $(\textsc{Encode},\textsc{Decode})$. Correctness says that one can “peel off” the noise that is added during the Encode operation and recover the original product of $A$ and $B$. The other (novel) feature, called transcript unpredictability, says that the only way to learn a piece of the transcript of the computation is to directly compute it. Satisfying each of the two properties separately is trivial, and the challenge is of course to have them hold at the same time.
@@
A transcript unpredictable encoding scheme $(\textsc{Encode},\allowbreak\textsc{Decode})$ satisfies the following properties:
@@
$\epsilon$-transcript unpredictability: There is a constant $C>0$ such that for every two $n\times n$ matrices $A,B$, and every $r$, every iterative numerical algorithm $\mathcal{A}$ that runs in time $t=t(n)$ cannot compute the transcript of the product of the noisy matrices $A^{\prime},B^{\prime}$, except with probability $C\cdot t/t_{\textsc{MatMul}}\cdot\epsilon$. That is,
$\Pr\left[\begin{aligned} &\forall i,j,k\in[n/r]\colon\\ &\quad C^{(k)}_{i,j}=Z^{(k)}_{i,j}\end{aligned}\,\middle|\,\begin{aligned} &A^{\prime},B^{\prime}=\textsc{Encode}_{r}(\sigma,A,B)\\ &\{C^{(k)}_{i,j}\}_{i,j,k\in[n/r]}=\mathsf{Tr}(\textsc{MatMul}_{r},A^{\prime},B^{\prime})\\ &\mathcal{\{}Z^{(k)}_{i,j}\}_{(i,j,k)\in[n/r]}=\mathcal{A}(A^{\prime},B^{\prime})\\ \end{aligned}\right]\leq\frac{C\cdot t(n)}{t_{\textsc{MatMul}}(n)}\cdot\epsilon(n),$
@@
We refer to Appendix C for a more general abstraction (that applies to any function, and not only matrix multiplication), termed unpredictable encodings, that we believe is of independent interest.
@@ 6.3 From Transcript Unpredictability to a PoUW
We now argue that any algorithm that runs in time strictly less than $t_{\textsc{MatMul}}(n)$, cannot convince the verifier to output 1, except with small probability. Suppose that there is a $P^{\star}$ that runs in time less than $t_{\textsc{MatMul}}(n)$. The claim follows directly form the $\epsilon$-transcript unpredictability property of the $(\textsc{Encode},\textsc{Decode})$ operations, concluding that with probability at most $\epsilon$ such a $P^{*}$ will fool the verifier, concluding the claim.
@@ 6.4 A First Instantiation
First Implementation of Encode and Decode $\textsc{Encode}_{r}(\sigma,A,B)$: $\quad\sslash\quad A,B\in{\mathbb{F}}_{q}^{n\times n}$ 1. Parse $\sigma=E,F$, where $E,F\in{\mathbb{F}}_{q}^{n\times n}$. 2. Output $A^{\prime}=A+E$ and $B^{\prime}=B+F$. $\textsc{Decode}_{r}(\sigma,C^{\prime})$: $\quad\sslash\quad C^{\prime}\in{\mathbb{F}}_{q}^{n\times n}$ 1. Parse $\sigma=E,F$, where $E,F\in{\mathbb{F}}_{q}^{n\times n}$. 2. Compute $C^{\prime\prime}=A\cdot F+E\cdot(B+F)$ (using two invocations of $\textsc{MatMul}_{r}$ and a matrix addition). 3. Output $C=C^{\prime}-C^{\prime\prime}$. Algorithm 6.3
@@
The obvious downside is that a malicious solver can skip the Decode operation altogether and thereby save two matrix multiplications. In other words, the honest solver does at least three times the work it should have done only for computing $\textsc{MatMul}_{r}$, i.e., the protocol’s overhead is $\alpha\geq 3$. On the other hand, there is a $\mathsf{Solve}^{*}$ that does essentially only $t_{\textsc{MatMul}}(n)$ work, concluding that the protocol is not useful.
@@ 6.5 A Second Instantiation
Second Implementation of Encode and Decode $\textsc{Encode}_{r}(\sigma,A,B)$: $\quad\sslash\quad\sigma\in\{0,1\}^{\lambda},\quad A,B\in{\mathbb{F}}_{q}^{n\times n}$ 1. Parse $\sigma=E_{L},E_{R},F_{L},F_{R}$ as four matrices, where $E_{L},F_{L}\in{\mathbb{F}}_{q}^{n\times r}$ and $E_{R},F_{R}\in{\mathbb{F}}_{q}^{r\times n}$. 2. Compute $E=E_{L}\cdot E_{R}\text{ and }F=F_{L}\cdot F_{R}$ (invoking MatMul twice). 3. Output $A^{\prime}=A+E$ and $B^{\prime}=B+F$. $\textsc{Decode}_{r}(\sigma,C^{\prime})$: $\quad\sslash\quad\sigma\in\{0,1\}^{\lambda},\quad C^{\prime}\in{\mathbb{F}}_{q}^{n\times n}$ 1. Parse $\sigma=E_{L},E_{R},F_{L},F_{R}$ as four matrices, where $E_{L},F_{L}\in{\mathbb{F}}_{q}^{n\times r}$ and $E_{R},F_{R}\in{\mathbb{F}}_{q}^{r\times n}$. 2. Compute $C^{\prime\prime}=(A\cdot F_{L})\cdot F_{R}+E_{L}\cdot(E_{R}\cdot(B+F_{L}\cdot F_{R}))$ (invoking MatMul 5 times). 3. Output $C=C^{\prime}-C^{\prime\prime}$. Algorithm 6.4
@@
Here, as opposed to all of our previous suggestions, the PoUW is useful. Indeed, the honest solver $\mathsf{Solve}$ is doing exactly one product of two $n\times n$ matrices, as all of the products in Encode and Decode are products of $n\times n$ and $n\times r$ matrices, that cost $O(n^{2}\cdot r)$. Overall, by a direct analysis, the overhead is
$\alpha(n)=O(n^{2}+n^{2}\cdot r)/t_{\textsc{MatMul}}(n).$
Note that $\alpha(n)=o(1)$ (unless fast matrix multiplication is used). By our assumption on the fastest way to compute the intermediate matrices (see Section 6.1, there is no algorithm that runs in time $o(t_{\textsc{MatMul}}(n))$, making our protocol a truly useful PoW.